# Lagrangian multiforms and dispersionless integrable systems

### v1

@@ Abstract
E.V.Ferapontov@lboro.ac.uk
M.Vermeeren@lboro.ac.uk
@@ 1 Lagrangian multiform formalism
We consider Lagrangian multiforms
@@
Here $L_{ij}$ are some functions of the jet variables $x^{i},x^{j},u,u_{i},u_{j},u_{ij}$ where $u$ is a function of the independent variables $x^{1},\dots,x^{n}$ and $u_{i}=u_{x^{i}},\ u_{ij}=u_{x^{i}x^{j}}$, etc, denote partial derivatives. We assume $L_{ji}=-L_{ij}$. For a surface $\Gamma$ in the space of independent variables, we consider the action integral $S_{\Gamma}=\int_{\Gamma}\mathcal{L}$. Lagrangian multiform theory requires that this action is critical for every choice of surface $\Gamma$, with respect to variations of $u$. (This requirement is also known as the pluri-Lagrangian principle.) This is the case if and only if the following three groups of multiform Euler-Lagrange equations hold:
@@
where $\partial_{i}=\frac{\partial}{\partial x^{i}}$, etc.
@@
Note that the definition of variational derivatives remains the same: the additional derivatives are only taken in the variables corresponding to the indices of $\delta$. In particular:
@@
For this framework to be nontrivial, we need to impose the condition that the multiform Euler-Lagrange equations (also known as multi-time Euler-Lagrange equations) are in involution, i.e. that the overdetermined system (1) can be solved, given suitable boundary conditions.
The fundamental idea of Lagrangian multiforms was established in [17]. The multiform Euler-Lagrange equations for a 2-form $\mathcal{L}$ depending on the second jet bundle first appeared in [31] and the multiform Euler-Lagrange equations for $\mathcal{L}$ depending on any jet bundle were derived in [32]. This derivation makes use of surfaces that are made up of flat pieces in coordinate directions. Then equation (1a) comes from the flat pieces, (1b) from the edges where two flat pieces meet, and (1c) from the corners where three or more flat pieces meet. For this reason, (1b) are called planar equations, (1b) edge equations and (1c) corner equations. Other derivations of the multiform Euler-Lagrange equations are possible. One can obtain equations (1) by restricting $\mathcal{L}$ to an arbitrary plane in the space of independent variables (not necessarily coordinate planes) and calculating the standard Euler-Lagrange equation, which should be satisfied for any such plane. Furthermore, in [30] it was pointed out that equivalent equations are obtained by taking pointwise variations of $\mathrm{d}\mathcal{L}$.
Lagrangian multiforms and conservation laws have been touched upon in [23, 29], where Lagrangian multiforms were constructed for systems with known variational symmetries. These papers use Noether’s theorem to obtain conservation laws from which a multiform is then constructed. In the present work, we take conservation laws as our starting point and investigate the additional structure provided by Lagrangian multiform theory.
@@
$B_{ijk}$ $=u_{ijk}-\frac{1}{2}\left(\frac{u_{ij}u_{ik}}{u_{i}}+\frac{u_{ij}u_{jk}}{u_{j}}+\frac{u_{ik}u_{jk}}{u_{i}k}\right),$
@@
$(x^{i}-x^{j})u_{k}u_{ij}+(x^{k}-x^{i})u_{j}u_{ik}+(x^{j}-x^{k})u_{i}u_{jk}=0,$ (9)
as well as (7), which remains unchanged. Note that translationally non-invariant Veronese web equation (9) has appeared in [15, 17].
@@
Factorisations of $\mathrm{d}\mathcal{L}$ were first used in a narrower sense in [23] and [29]. Since then, arguments involving double zeros have inspired the construction of Lagrangian multiforms in several different settings [28, 2, 19, 26].
The Lagrangian multiform (4) has two important properties, which motivate the structure of the paper:
  • The relation d=0(mod $\mathrm{d}\mathcal{L}=0\ (\text{mod }A_{ijk}=0)$ suggests that $\mathcal{L}$ is a (second-order) conservation law of the Veronese web hierarchy (6), for which $B_{ijk}$ can be viewed as its characteristic. In Section 2 we provide further examples of this kind by calculating second-order conservation laws of some linearly degenerate PDEs in 3D. It is yet to be explored to what extent conservation laws of integrable PDEs/hierarchies can be viewed as a common source of Lagrangian multiforms.
  • It will be shown in Section 3 that equations (6)-(7) coincide with Gibbons-Tsarev equations governing hydrodynamic reductions of the 4D second heavenly equation. This suggests that Gibbons-Tsarev equations in 4D constitute yet another natural source of Lagrangian multiforms.
@@ 2 Lagrangian multiforms and conservation laws of linearly degenerate PDEs
The examples below support a point of view of Lagrangian multiforms as (higher) conservation laws of integrable PDEs (hierarchies).
@@ 2.1 Veronese web hierarchy
in the first of these expressions, the index $i\in\{1,\dots,n\}$ is fixed, while $j$ varies. Taking their linear combination gives Lagrangian multiform
@@
which can be seen as a multi-parameter deformation of the Lagrangian multiform (4). The exterior derivative factorises as in equation (2), with
@@
$B_{ijk}$ $=u_{ijk}-\frac{1}{2}\left(\frac{u_{ij}u_{ik}}{u_{i}}+\frac{u_{ij}u_{jk}}{u_{j}}+\frac{u_{ik}u_{jk}}{u_{k}}\right)$
$\quad+\frac{n^{i}u_{j}u_{k}}{2(c^{i}-c^{j})(c^{i}-c^{k})u_{i}}+\frac{n^{j}u_{i}u_{k}}{2(c^{j}-c^{i})(c^{j}-c^{k})u_{j}}+\frac{n^{k}u_{i}u_{j}}{2(c^{k}-c^{i})(c^{k}-c^{j})u_{k}}-\varepsilon\frac{u_{i}u_{j}u_{k}}{2}.$
Hence, by Theorem 1, the system $A_{ijk}=0$, $B_{ijk}=0$ implies the multiform Euler-Lagrange equations, which are (compare (5)):
@@ 2.2 Translationally non-invariant Veronese web hierarchy
respectively.
@@ 2.3 Mikhalev equation
It first appeared in [18], in the context of the Hamiltonian formalism of KdV type hierarchies. It possesses three second-order conservation laws,
@@
For each of them, the differential $\mathrm{d}\mathcal{L}=(F_{1}+G_{2}+H_{3})\,\mathrm{d}x^{1}\wedge\mathrm{d}x^{2}\wedge\mathrm{d}x^{3}$ factorises as
@@
where the characteristic $\Sigma$ is a differential expression in $u$ such that the combined system $\{u_{33}-u_{12}+u_{3}u_{11}-u_{1}u_{13}=0,\ \Sigma=0\}$ is involutive. Thus, these conservation laws give rise to Lagrangian multiforms, where $L_{12}=H$, $L_{13}=-G$, and $L_{23}=F$.
@@
$\frac{\delta_{31}G}{\delta u_{1}}+\frac{\delta_{23}F}{\delta u_{2}}=-\partial_{1}\frac{\partial G}{\partial u_{11}}-0=2u_{111}.$
@@
the deformed ~$\tilde{\mathcal{L}}=\tilde{F}\,\mathrm{d}x^{2}\wedge\mathrm{d}x^{3}+\tilde{G}\,\mathrm{d}x^{3}\wedge\mathrm{d}x^{1}+\tilde{H}\,\mathrm{d}x^{1}\wedge\mathrm{d}x^{2}$ is also a Lagrangian multiform. Note that the second equation is the potential KdV equation.
@@
$\qquad+4u_{1}u_{13}u_{33}-u_{33}^{2}.$
@@
Similar second-order conservation laws, as well as the associated Lagrangian multiforms, can be constructed for other 3D linearly degenerate second-order dispersionless integrable PDE as classified in [6].
@@ 3 Lagrangian multiforms and Gibbons-Tsarev equations
In the most general setting, the method of hydrodynamic reductions (that leads to equations of Gibbons-Tsarev type) applies to quasilinear PDEs of the form
@@
or any other classes of PDEs transformable into form (13), see below. Here $u$ is a (column) vector-function of $d$ independent variables $x^{1},\dots,x^{d}$, and $u_{x^{i}}$ denote partial derivatives (the matrices $A^{i}(u)$ do not need to be square). For definiteness, let us consider $4$-dimensional PDEs ($d=4$) with four independent variables $t,x,y,z$. Let us look for exact solutions of (13) in the form ${u}={u}(R)$ where the Riemann invariants $R=\{R^{1},...,R^{n}\}$ solve a triple of commuting diagonal systems
@@
$i\neq j,\ \lambda_{j}^{i}=\frac{\partial}{\partial{R^{j}}}\lambda^{i}$, etc. Substituting ${u}(R)$ into (13) and using (14), one arrives at an over-determined system of equations for $u(R)$ and the characteristic speeds $\lambda^{i}(R),\mu^{i}(R),\eta^{i}(R)$, known as Gibbons-Tsarev equations. These equations imply, in particular, that the characteristic speeds $\lambda^{i},\mu^{i}$ and $\eta^{i}$ must satisfy an algebraic relation which can be interpreted as the dispersion relation of system (13).
One can show that the maximum “amount” of $n$-component reductions a $d$-dimensional PDE may possess is parametrised, modulo changes of variables $R^{i}\to f^{i}(R^{i})$, by $(d-2)n$ arbitrary functions of one variable [3, 5].
A $d$-dimensional PDE is said to be integrable if its $n$-component reductions are locally parametrised by $(d-2)n$ arbitrary functions of one variable [3], [5]. Our main observation is that Gibbons-Tsarev type systems governing such reductions possess a Lagrangian multiform representation, in all dimensions $d\geq 4$. Below we provide Lagrangian multiform representations of Gibbons-Tsarev equations for some well-known integrable PDEs.
@@ 3.1 Second heavenly equation
The so-called Plebański second heavenly equation,
@@
    It follows that there exists some function $P$ such that $P_{i}=\frac{\partial}{\partial R^{i}}P$ for all $i$.
@@
    $-u_{ij}+2\frac{\mu^{j}-\mu^{i}}{c^{j}-c^{i}}u_{i}u_{j}=0.$
  • (1b) with respect to $\mu^{i}$ instead of $u$ yields the trivial equation $-u_{i}=-u_{i}$.

### v2

@@ Abstract
This version of the article has been accepted for publication, after peer review but is not the Version of Record and does not reflect post-acceptance improvements, or any corrections. The Version of Record is available online at: https://doi.org/10.1007/s11005-025-02016-w
@@ 1 Lagrangian multiform formalism
Lagrangian multiform theory provides a variational principle for systems that are integrable in the sense of multidimensional consistency. It applies in the discrete case (where multidimensional consistency takes the form of commuting maps, consistency around the cube, etc), in the continuous case (where multidimensional consistency takes the form of commuting flows or involutive PDEs), and in semi-discrete settings.
The central object in continuous Lagrangian multiform theory is a differential $d$-form. Typically, a Lagrangian $d$-form describes a system of equations in $d$ independent variables each. For example,
  • a continuous Lagrangian 1-form describes commuting ODEs and can often be associated to a system of Poisson-commuting Hamiltonians [43, 39];
  • a continuous Lagrangian 2-form describes a hierarchy of 2D PDEs, such as the (potential) Korteweg-de Vries hierarchy [40, 41];
  • continuous Lagrangian 3-forms describe hierarchies of 3D PDEs, such as the KP hierarchy [38, 25].
This does not mean that all the Euler-Lagrange equations of a Lagrangian $d$-form are $d$-dimensional PDEs, but rather that the system of Euler-Lagrange equations is of a $d$-dimensional nature. In this paper we will show examples of 2-forms for which the variational equations appear as a system of PDEs in 3 variables each. In addition, this paper establishes a new connection between Lagrangian multiforms and conservation laws, and provides examples of Lagrangian multiforms for Gibbons-Tsarev type systems.
1.1 Multiform Euler-Lagrange equations
We consider Lagrangian 2-forms
@@ 1.1 Multiform Euler-Lagrange equations
Here $L_{ij}$ are some functions of the jet bundle variables $x^{i},x^{j},u,u_{i},u_{j},u_{ij}$ where $u$ is a function of the independent variables $x^{1},\dots,x^{n}$ and $u_{i}=u_{x^{i}},\ u_{ij}=u_{x^{i}x^{j}}$, etc, denote partial derivatives. We assume $L_{ji}=-L_{ij}$. For a two-dimensional surface $\Gamma$ in the space of independent variables, we consider the action integral $S_{\Gamma}=\int_{\Gamma}\mathcal{L}$. Lagrangian multiform theory requires that this action is critical for every choice of surface $\Gamma$, with respect to variations of $u$. (This requirement is also known as the pluri-Lagrangian principle [4, 40].) This is the case if and only if the following three groups of multiform Euler-Lagrange equations hold:
@@
where $\partial_{i}=\frac{\partial}{\partial x^{i}}$, etc. There is no summation over repeated indices.
@@
Note that the definition of variational derivatives remains the same: the additional derivatives are only taken in the variables corresponding to the indices of $\delta$. In particular,
@@
For this framework to be nontrivial, we need to impose the condition that the multiform Euler-Lagrange equations (also known as multi-time Euler-Lagrange equations) are in involution, i.e. that the overdetermined system (1) admits solutions for suitable generic boundary conditions.
The fundamental idea of Lagrangian multiforms was established in [22]. The multiform Euler-Lagrange equations for a 2-form $\mathcal{L}$ depending on the second jet bundle first appeared in [40] and the multiform Euler-Lagrange equations for $\mathcal{L}$ depending on any jet bundle were derived in [41]. This derivation makes use of surfaces that are made up of flat pieces in coordinate directions. Then equation (1a) comes from the flat pieces, (1b) from the edges where two flat pieces meet, and (1c) from the corners where three or more flat pieces meet. For this reason, (1a) are called planar equations, (1b) edge equations and (1c) corner equations. Other derivations of the multiform Euler-Lagrange equations are possible. One can obtain equations (1) by restricting $\mathcal{L}$ to an arbitrary plane in the space of independent variables (not necessarily coordinate planes) and calculating the standard Euler-Lagrange equation, which should be satisfied for any such plane. Furthermore, in [38] it was pointed out that equivalent equations are obtained by taking pointwise variations of $\mathrm{d}\mathcal{L}$.
The connection between Lagrangian multiforms and conservation laws has been touched upon in [40, 29] and was used in [30, 37] to construct Lagrangian 2-forms for systems with known variational symmetries. These papers start with a known Lagrangian and use Noether’s theorem to obtain conservation laws from which a multiform is then constructed. In the present work, we take conservation laws as our starting point and investigate the additional structure provided by Lagrangian multiform theory.
1.2 Double zero property
@@
$B_{ijk}$ $=2u_{ijk}-\left(\frac{u_{ij}u_{ik}}{u_{i}}+\frac{u_{ij}u_{jk}}{u_{j}}+\frac{u_{ik}u_{jk}}{u_{k}}\right),$
@@
$(x^{i}-x^{j})u_{k}u_{ij}+(x^{k}-x^{i})u_{j}u_{ik}+(x^{j}-x^{k})u_{i}u_{jk}=0$ (9)
and (7), which remains unchanged. Note that translationally non-invariant Veronese web equation (9) has appeared in [20, 22].
@@
Factorisations of $\mathrm{d}\mathcal{L}$ were first used in a narrower sense in [30] and [37]. Since then, arguments involving double zeros have inspired the construction of Lagrangian multiforms in several different settings [36, 6, 25, 33, 34].
The Lagrangian multiform (4) has two important properties that motivated the structure of the present paper:
  • The relation $\mathrm{d}\mathcal{L}=0\ (\text{mod }A_{ijk}=0)$ suggests that $\mathcal{L}$ is a (second-order) conservation law of the Veronese web hierarchy (6), for which $B_{ijk}$ can be viewed as its characteristic. In Section 2 we provide further examples of this kind by using second-order conservation laws of some linearly degenerate PDEs in 3D such as the Veronese web hierarchy and the Mikhalev equation. It is yet to be explored to what extent conservation laws of integrable PDEs/hierarchies can be viewed as a common source of Lagrangian multiforms.
  • It will be shown in Section 3 that equations (6)–(7) coincide with the Gibbons-Tsarev equations governing hydrodynamic reductions of the 4D second heavenly equation. Further examples (such as the first heavenly equation and a 6D version of the second heavenly equation) suggest that Gibbons-Tsarev type equations governing hydrodynamic reductions of linearly degenerate dispersionless integrable PDEs possess a Lagrangian multiform representation. Such Gibbons-Tsarev equations are only of interest in dimensions $d\geq 4$ as, according to [26], Gibbons-Tsarev equations of three-dimensional linearly degenerate PDEs are linearisable. In contrast, Gibbons-Tsarev equations of higher-dimensional linearly degenerate PDEs are not linearisable (not even Darboux integrable).
This paper is of an experimental nature, in the sense that it presents a number of intriguing examples. They support two observations that relate to the points above and appear to be new: (1) that Lagrangian 2-forms appear as (higher) conservation laws of various 3D integrable PDEs; (2) that Lagrangian multiforms appear in the context of Gibbons-Tsarev equations governing hydrodynamic reductions of various heavenly-type PDEs in 4D.
@@ 2 Lagrangian multiforms and conservation laws of linearly degenerate PDEs
The examples below support a point of view of Lagrangian multiforms as (higher) conservation laws of integrable PDEs (hierarchies). Each of these examples starts from a 3D PDE with a known conservation law, without assuming any Lagrangian structure a priori. This is in contrast to previous works [29, 30, 37], where Lagrangian multiforms are constructed by starting from a known Lagrangian and its variational symmetries.
@@ 2.1 Veronese web hierarchy
in the first of these expressions, the index $i\in\{1,\dots,n\}$ is fixed, while $j$ varies (here under the ‘order’ of a conservation law we understand the highest order of partial derivatives of $u$ involved). Taking their linear combination gives Lagrangian multiform
@@
which can be seen as a multi-parameter deformation of the Lagrangian multiform (4); here the constants $n^{i}$ and $\varepsilon$ are the deformation parameters. The exterior derivative factorises as in equation (2), with
@@
$B_{ijk}$ $=2u_{ijk}-\left(\frac{u_{ij}u_{ik}}{u_{i}}+\frac{u_{ij}u_{jk}}{u_{j}}+\frac{u_{ik}u_{jk}}{u_{k}}\right)$
$\quad+\frac{n^{i}u_{j}u_{k}}{(c^{i}-c^{j})(c^{i}-c^{k})u_{i}}+\frac{n^{j}u_{i}u_{k}}{(c^{j}-c^{i})(c^{j}-c^{k})u_{j}}+\frac{n^{k}u_{i}u_{j}}{(c^{k}-c^{i})(c^{k}-c^{j})u_{k}}-\varepsilon u_{i}u_{j}u_{k}.$
Hence, by Theorem 1, the system $A_{ijk}=0$, $B_{ijk}=0$ implies the multiform Euler-Lagrange equations, which are (compare with (5)):
@@
We emphasise that the system $A_{ijk}=0$, $B_{ijk}=0$ is involutive, with the general solution depending on $2n$ arbitrary functions of one variable.
@@ 2.2 Translationally non-invariant Veronese web hierarchy
respectively. As in the previous case, the system $A_{ijk}=0$, $B_{ijk}=0$ is involutive, with the general solution depending on $2n$ arbitrary functions of one variable.
@@ 2.3 Mikhalev equation
It first appeared in [23], in the context of the Hamiltonian formalism of KdV type hierarchies. It possesses four first-order and seven second-order conservation laws (see [1] for the characteristics of the latter),
@@
For any such conservation law (or a linear combination thereof), the differential $\mathrm{d}\mathcal{L}=(F_{1}+G_{2}+H_{3})\,\mathrm{d}x^{1}\wedge\mathrm{d}x^{2}\wedge\mathrm{d}x^{3}$ factorises as
@@
where the characteristic $\Sigma$ is a differential expression in $u$ such that the combined system $\{u_{33}-u_{12}+u_{3}u_{11}-u_{1}u_{13}=0,\ \Sigma=0\}$ is involutive. Thus, these conservation laws give rise to Lagrangian multiforms, where $L_{12}=H$, $L_{13}=-G$, and $L_{23}=F$. Below we discuss some of the simplest examples of this constuction.
@@
$\frac{\delta_{31}G}{\delta u_{1}}+\frac{\delta_{23}F}{\delta u_{2}}=-\partial_{1}\frac{\partial G}{\partial u_{11}}+0=2u_{111}.$
@@
the deformed $\tilde{\mathcal{L}}=\tilde{F}\,\mathrm{d}x^{2}\wedge\mathrm{d}x^{3}+\tilde{G}\,\mathrm{d}x^{3}\wedge\mathrm{d}x^{1}+\tilde{H}\,\mathrm{d}x^{1}\wedge\mathrm{d}x^{2}$ is also a Lagrangian multiform. Note that the second equation is the potential KdV equation.
@@
$\qquad+4u_{1}u_{13}u_{33}-u_{33}^{2},$
$\tilde{G}$ $=-\left(u_{1}u_{11}-u_{13}\right)^{2},$
$\tilde{H}$ $=-u_{1}^{3}u_{11}^{2}+2u_{1}^{2}u_{11}u_{13}+2u_{1}u_{11}^{2}u_{3}-2u_{1}u_{11}u_{12}-u_{1}u_{13}^{2}-2u_{11}u_{13}u_{3}+2\,u_{12}u_{13}.$
@@
Similar second-order conservation laws, as well as the associated Lagrangian multiforms, can be constructed for other 3D linearly degenerate second-order dispersionless integrable PDE as classified in [11]. We refer to [1, 24] where the characteristics of these conservation laws were calculated using third-order symmetries of the cotangent coverings of these equations.
@@ 3 Lagrangian multiforms and Gibbons-Tsarev equations
In this section we present Lagrangian multiforms for some Gibbons-Tsarev equations. To introduce Gibbons-Tsarev equations, we begin with a brief review of the method of hydrodynamic reductions. In the most general setting, the method of hydrodynamic reductions applies to quasilinear PDEs of the form
@@
or any other classes of PDEs transformable into form (13), see below. Here $u$ is a (column) vector-function of $d$ independent variables $x^{1},\dots,x^{d}$, and $u_{x^{i}}$ denote partial derivatives (the matrices $A^{i}(u)$ do not need to be square). For definiteness, let us consider $4$-dimensional PDEs ($d=4$) with four independent variables $t,x,y,z$. Let us look for solutions of (13) in the form u=u(R){u}={u}(R) where the Riemann invariants $R=\{R^{1},...,R^{n}\}$ solve a triple of commuting diagonal systems
@@
$i\neq j,\ \lambda_{j}^{i}=\frac{\partial}{\partial{R^{j}}}\lambda^{i}$, etc. Substituting u(R){u}(R) into (13) and using (14), one arrives at an over-determined system of equations for $u(R)$ and the characteristic speeds $\lambda^{i}(R),\mu^{i}(R),\eta^{i}(R)$, known as Gibbons-Tsarev equations. These equations imply, in particular, that the characteristic speeds $\lambda^{i},\mu^{i}$ and $\eta^{i}$ must satisfy an algebraic relation which can be interpreted as the dispersion relation of system (13).
One can show that the maximum “amount” of $n$-component reductions a $d$-dimensional PDE may possess is parametrised, modulo changes of variables $R^{i}\to f^{i}(R^{i})$, by (d2)n(d-2)n arbitrary functions of one variable.
Definition ([7, 9]).
A $d$-dimensional system (13) is said to be integrable if its $n$-component reductions are locally parametrised by (d2)n(d-2)n arbitrary functions of one variable.
Although integrability in the sense of hydrodynamic reductions is somewhat different from the familiar solitonic integrability, is has all attributes of any “reasonable” definition of integrability:
  • it is based on exact solutions (coming from the method of hydrodynamic reductions) which are locally dense in the space of all solutions of a given PDE; these solutions can be considered as natural ‘dispersionless analogues’ of multisoliton solutions of conventional integrable PDEs;
  • it is algorithmically verifiable;
  • it leads to classification results of integrable PDEs within various particularly interesting classes;
  • it can be reformulated geometrically as a certain involutivity condition of the principal symbol of the given PDE system.
All these properties are thoroughly discussed in [7, 9, 10, 2, 3]. Our main observation is that many Gibbons-Tsarev type equations governing hydrodynamic reductions of various dispersionless integrable systems possess a Lagrangian multiform representation, in all dimensions $d\geq 4$. Below we provide Lagrangian multiform representations of Gibbons-Tsarev equations for some well-known integrable PDEs.
@@ 3.1 Second heavenly equation
Plebański’s second heavenly equation,
@@
    It follows that there exists some function $P$ such that $P_{i}=\frac{\partial}{\partial R^{i}}P$ for all ii.
@@
    $u_{ij}+2\frac{\mu^{j}-\mu^{i}}{c^{j}-c^{i}}u_{i}u_{j}=0.$
  • (1b) with respect to $\mu^{i}$ instead of $u$ yields the trivial equation $u_{i}=-u_{i}$.